An office worker smacks a table, bringing his hand to rest in 0.0025 seconds from an initial speed of 4.00 m/s. How much force is exerted on the table, taking the effective mass of his forearm and hand to be 1.5 kg?
1
Expert's answer
2020-11-16T07:50:42-0500

The force is (according to Newton's second law):


"F=ma."

The acceleration is how quickly speed changes:


"a=\frac{v_f-v_i}{t}."

Thus, the force is


"F=m\frac{v_f-v_i}{t}=1.5\frac{0-4}{0.0025}=2400\text{ N}."